# Architectural Patterns for Designing Quantum Artificial Intelligence Systems

### v2

@@ 8 Conclusions
The unique software architectural solutions concerning quantum AI primarily revolve around determining which tasks to delegate to a quantum computer and how quantum and classical software components communicate with each other. During the systematic mapping study, we identified ten architectural patterns in quantum AI systems: seven related to the quantum-classic split and three related to quantum middleware. These patterns constitute the core of the pattern catalog developed in this work, which aims to guide software architects in effectively utilising quantum components in their projects. Each architectural pattern involves trade-offs in achieving the system quality attributes, such as efficiency, scalability, trainability, simplicity, portability, and deployability. The most commonly used patterns for small-scale proof-of-concept applications are the quantum monolith and quantum feature engineering. When attempting to scale the system to handle larger volumes of information within the constraints of NISQ-era hardware, the corresponding challenges are typically addressed using architectural patterns such as the quantum head or quantum intermediate layer. In these architectures, part of the inference task, as well as the reduction of data dimensionality, is delegated to classical components. The primary approach for organising communication between classical and quantum components is through the use of a service-oriented architecture. In quantum AI systems, quantum components can be integrated into trainable inference models or function as accelerators for specific tasks.

### v3

@@ 8 Conclusions
The unique software architectural solutions concerning quantum AI primarily revolve around determining which tasks to delegate to a quantum computer and how quantum and classical software components communicate with each other. During the systematic mapping study, we identified ten architectural patterns in quantum AI systems: seven related to the quantum-classic split and three related to quantum middleware. These patterns constitute the core of the pattern catalog developed in this work, which aims to guide software architects in effectively utilising quantum components in their projects. Each architectural pattern involves trade-offs in achieving the system quality attributes, such as efficiency, scalability, trainability, simplicity, portability, and deployability. The most commonly used patterns for small-scale proof-of-concept applications are the quantum monolith and quantum feature engineering. When attempting to scale the system up to handle larger volumes of information within the constraints of NISQ-era hardware, the corresponding challenges are typically addressed using architectural patterns such as the quantum head or quantum intermediate layer. In these architectures, part of the inference task, as well as the reduction of data dimensionality, is delegated to classical components. The primary approach for organising communication between classical and quantum components is through the use of a service-oriented architecture and service wrappers. In quantum AI systems, quantum components can be integrated into trainable inference models or function as accelerators for specific tasks.
The architectural patterns identified in this study describe architectural solutions that are specifically tailored for quantum AI systems. These patterns are crucial for addressing key issues such as computational efficiency, scalability and data flow management in those systems. By adopting these architectural patterns, developers can mitigate risks, improve maintainability, and optimize performance in quantum AI systems.